Steves adverage score in biology after 8 tests is a 79. In order to get an 80 what is the minimum score e needs to get on the final test. Do not round up

A. 95
B. 88
C. 85
D. 82

Answer by amalm06(224)

(a+b+c+d+e+f+g+h)/8=79
So a+b+c+d+e+f+g+h=632
In order to get the average score to 80, include 1 more test, call it i
(a+b+c+d+e+f+g+h+i)/9=80
Since a+b+c+d+e+f+g+h=632, we have
(632+i)/9=80
632+i=720
i=88 (Answer)